Fermi National Laboratory


Accelerator Update

The integrated luminosity for last week was 12.30 pb-1.

Monday February 9
The day shift began with Operations monitoring store 3224 and with Antiproton Source (Pbar) experts conducting reverse proton studies.

At 10:23 AM, operators made Booster ready for access. A Linac expert accessed the I- Source to do maintenance. Technicians also accessed MiniBooNE. Linac techs worked on two RF stations.

Operations established beam to Booster at 8 PM.

Operations established beam to MiniBooNE at 8:14 PM.

Pbar began stacking at 8:15 PM.

Earlier in the evening, the Collider Detector at Fermilab (CDF) had a cryo system problem that had held them off until it was repaired at 8:30 PM.

At 11 PM, operators reported that the I- Source had sparked often throughout the evening shift.

Tuesday February 10
The midnight shift began with Operations monitoring store 3224, with Pbar stacking, and with MiniBooNE taking beam.

Operations reported at 2:44 AM that the I- Source was tripping off regularly. An operator lowered the command voltage slightly, which helped.

At 5:06 AM, a problem developed with the Tevatron (TeV) LCW pump located at the sector B4 service building. An expert told the Duty Electrician to bypass the pump. He thought the temperatures and pressures would hold, but that Operations should monitor the system closely.

An operator reported at 7:14 AM that a Booster RF station (BRF13) tripped off about twenty times during the midnight shift.

At 9:58 AM, a Main Injector (MI) expert began a slip stacking study.

Operations terminated store 3224 at 11:36 AM.

At 12:54 PM, Operations put the TeV into shot setup.

At 3:21 PM, Operations established store 3226 with an initial luminosity of 51.38E30. Pbar resumed stacking.

Operations established beam to Switchyard for training at 4:23 PM, but then turned it off.

Operations and experts from Recycler, Pbar, and MI began transferring shots of antiprotons from Pbar to the Recycler at 4:26 PM. These transfers ended around 5:38 PM. Experts resumed stacking and Switchyard studies.

Switchyard studies ended at 7:41 PM.

At 11:33 PM, MI suffered a vacuum burst in its RF sector. Everything reset okay.

Wednesday February 11
The midnight shift began with Operations monitoring store 3226, with Pbar stacking, with MiniBooNE taking beam, and with Recycler maintaining a small stash.

At 4:23 AM, operators reported that the I- Source had been tripping repeatedly. They lowered the voltage.

Today's Plans
The plans for today are to continue with stacking and monitoring the store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/11/2004   email Fermilab